OpenStudy (anonymous):

can any one solve this?(log(a)(x^3))+(log(a)(sqrtx))= answer has to be in klog(a)x form thanks already :)

OpenStudy (amistre64):

is that base a?

OpenStudy (amistre64):

log[a](x^3) like this?

OpenStudy (amistre64):

3loga(x) + (1/2)loga(x)

OpenStudy (amistre64):

loga(x)(3+(1/2)) ??

OpenStudy (anonymous):

Amistre, I think contact has been lost lol.

OpenStudy (amistre64):

looks right lol (6/2) loga (x)

OpenStudy (amistre64):

:) but the math remains lol

OpenStudy (anonymous):

Hahahaa, so very true!

OpenStudy (amistre64):

(7/2) loga(x)

OpenStudy (anonymous):

yes a is the base, soz my internet played up :S

OpenStudy (amistre64):

can we factor out the loga(x) like that tho?

OpenStudy (anonymous):

\[\log_{a} x^3+\log_{a} sqrtx\]

OpenStudy (amistre64):

if we multiply it togeter again we get: loga[sqrt(x^7)] so its good

OpenStudy (amistre64):

(7/2) loga(x) is your answer

OpenStudy (anonymous):

where did 7 come from? soz but i dont get the stuff you lot did at the top :S

OpenStudy (amistre64):

3 + 1/2 = 3' 1/2 to get it into an improper fraction (topheavy) we multiply 3 by the denom and add the numerator; 3(2) = 6 + 1 = 7.... 7/2

OpenStudy (amistre64):

OR we could go the route of changing addition of logs back into multiplication of logs

OpenStudy (anonymous):

ok :) multiplication logs?

OpenStudy (amistre64):

loga(x^3 sqrt(x)) = loga(sqrt(x^7)) = loga(x^(7/2)) (7/2) loga(x)

OpenStudy (amistre64):

loga(x) + loga(y) = loga(xy)

OpenStudy (amistre64):

loga(x^n) = n loga(x) as well

OpenStudy (anonymous):

ok so 7/2logax is the answer right?

OpenStudy (amistre64):

yes; (7/2) loga(x) is the answer

OpenStudy (anonymous):

\[7/2\log_{a} x\]

OpenStudy (anonymous):

yay! ill give u all medals :) thx

OpenStudy (amistre64):

im gonna need a new cigar box to keep these things in ;)

OpenStudy (anonymous):

:p